OpenStudy (anonymous):

limit as x goes to 1+ (x^2-1)^(x-1) using l'hopitals rule

OpenStudy (turingtest):

I no of no other way to do this than to exponentiate this first

OpenStudy (turingtest):

\[\lim_{x\to a}f(x)=e^{\lim_{x\to a}\ln f(x)}=\exp\left(\lim_{x\to a}\ln f(x)\right)\]do you know this trick?

OpenStudy (anonymous):

so (x-1)ln(x^2-1)

OpenStudy (anonymous):

\[(x-1)\ln(x^2-1)\]

OpenStudy (turingtest):

yes, and now I would do the 1/t sub I bet there is a better way though... that anonymous is about to show us

OpenStudy (anonymous):

ok lol ill wait for his response

OpenStudy (turingtest):

oh no, my sub would be wrong

OpenStudy (turingtest):

maybe let x-1=1/t and let t go to infinity

OpenStudy (turingtest):

\[(x-1)[\ln(x-1)+\ln(x+1)]={[\ln(x-1)+\ln(x+1)]\over(x-1)^{-1}}\]I am very stuck here

OpenStudy (turingtest):

Zarkon to the rescue!

OpenStudy (zarkon):

do you know that \[\lim_{x\to 0^+}x^x=1\]?

OpenStudy (turingtest):

ooooooh

OpenStudy (turingtest):

\[\lim_{x\to1^+}(x-1)^{x-1}(x+1)^{x-1}=\lim_{x\to1^+}(x+1)^{x-1}\]that is a step in the right direction I suppose

OpenStudy (zarkon):

yep...and the answer follows

OpenStudy (zarkon):

\[2^0=1\]

OpenStudy (turingtest):

oh yeah, lol it's defined now

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1^+}(x^2-1)^{(x-1)}=\lim_{x \rightarrow 1^+}(x^2-1)^x/(x^2-1)\] lim (x^2-1)(x^2-1)(x^2-1)......(x^2-1)/(x^2-1) x-> 1^+ ^ x times lim (x^2-1)(x^2-1).....(x^2-1) x->1^+ ^ x-1 times =(1-1)(1-1).....(1-1)=0

OpenStudy (turingtest):

...but the limit is 1

OpenStudy (anonymous):

isn't 0^0 undefined?

OpenStudy (anonymous):

sorry to ask

OpenStudy (zarkon):

\[0^0\] is undefined...but \[\lim_{x\to 0^+}x^x\] is not

OpenStudy (zarkon):

not undefined that is ;)

OpenStudy (zarkon):

remember in the limit x is never 0

OpenStudy (turingtest):

\[\lim_{x\to0^+}x^x=\lim_{x\to0^+}\exp(x\ln x)=\exp(\lim_{t\to\infty}{-\ln t\over t})=\exp(\lim_{t\to\infty}-\frac1t)\]\[=\exp(0)=e^0=1\]as @Zarkon said

OpenStudy (anonymous):

oh ok I forgot product rule

OpenStudy (anonymous):

thanks everyone wish you guys could all be my study buddies! :D

OpenStudy (turingtest):

we are ;) welcome!

OpenStudy (anonymous):

i mean like in person study buddies lol!